# NaturalReasoning: Reasoning in the Wild with 2.8M Challenging Questions

### v3

@@ 9 Conclusion
Although our study already validates NaturalReasoning ’s value for large-scale offline training—covering supervised distillation and preference-based self-training (RFT, DPO)—we did not pursue complementary directions such as online RL with reference-answer rewards or an exhaustive survey of reward models; these remain natural extensions for future work. This paper seeks to improve reasoning capabilities of large language models through leveraging pretraining corpora. While our efforts are focused on curating high-quality, diverse data, models trained using this data may exhibit undesirable behavior not examined in our work. Therefore, comprehensive evaluation would be needed to evaluate and address any potential pre-existing or existing biases in LLMs which leverage this data.
@@ Appendix B Example questions
Example questions with single answer, short answer, long answer are shown in Table 4.
@@
[…]
@@
(2) Prove that the relation $R=\{(1,2),(1,3),(1,4),(2,1),(2,3),(2,4),(3,1),(3,2),(3,4),(4,1),(4,1),$ (4,3)}(4,3)\}( 4 , 3 ) } is not transitive. Use the definition of transitivity to show that there exist elements $x_{0},y_{0},z_{0}$ such that $(x_{0},y_{0})\in R$ and $(y_{0},z_{0})\in R$ but $(x_{0},z_{0})\notin R$.
@@ C.1 Generation
We use vllm for all generations. For annotating documents and synthesizing questions, we use greedy decoding (i.e. temperature=00). For response generation for each question in NaturalReasoning, we use temperature=$0.7$ top_p=$0.9$. Responses used in unsupervised self-training experiments are sampled using temperature={0, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0, 1.2} to encourage response diversity.
@@ Appendix F Reference Answer Usefulness
Unfiltered SFT 12349 31.82 49.92 40.87
Filtered SFT 7646 32.15 50.06 41.11
Appendix G Prompts

### v4

@@ 9 Conclusion
Although our study already validates NaturalReasoning ’s value for large-scale offline training—covering supervised distillation and preference-based self-training (RFT, DPO)—we also conduct preliminary experiments using online RL with verifiable rewards with General Verifier, which show promising gains even with limited training. A more systematic exploration of reinforcement learning paradigms, including alternative reward models and scaling strategies remains natural extensions for future work. This paper seeks to improve reasoning capabilities of large language models through leveraging pretraining corpora. While our efforts are focused on curating high-quality, diverse data, models trained using this data may exhibit undesirable behavior not examined in our work. Therefore, comprehensive evaluation would be needed to evaluate and address any potential pre-existing or existing biases in LLMs which leverage this data.
@@ Appendix B Example questions
Example questions with single word answer, short answer, long answer are shown in Table 4.
@@
(4) Evaluate the definite integral $\int_{0}^{\pi}\cos^{2}(x)\sin^{7}(x)dx$ using an appropriate substitution method and provide a step-by-step solution.
@@
(2) Prove that the relation $R=\{(1,2),(1,3),(1,4),(2,1),(2,3),(2,4),(3,1),(3,2),(3,4),(4,1),(4,1),$ (4,3)}(4,3)\} is not transitive. Use the definition of transitivity to show that there exist elements $x_{0},y_{0},z_{0}$ such that $(x_{0},y_{0})\in R$ and $(y_{0},z_{0})\in R$ but $(x_{0},z_{0})\notin R$.
@@ C.1 Generation
We use vllm for all generations. For annotating documents and synthesizing questions, we use greedy decoding (i.e. temperature=0). For response generation for each question in NaturalReasoning, we use temperature=$0.7$ top_p=$0.9$. Responses used in unsupervised self-training experiments are sampled using temperature={0, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0, 1.2} to encourage response diversity.
@@ F.1 Data Filtering In Knowledge Distillation
Unfiltered SFT 12,349 31.82 49.92 40.87
Filtered SFT 7,646 32.15 50.06 41.11
F.2 Reinforcement Learning With Verifiable Rewards
We conduct preliminary experiments applying reinforcement with verifiable rewards (RLVR) on the NaturalReasoning dataset. Specifically, we sample a subset of questions whose reference answers are shorter than 10 words and train Llama3.1-8B-Instuct using GRPO [shao2024deepseekmathpushinglimitsmathematical] with the General Verifier [general-reasoner] as the reward model. Training is performed with a batch size of 768. Despite only 50 optimization steps, the model already exhibit noticeable performance gains over the untrained baseline across multiple reasoning benchmarks as shown in Table 7.
Table 7: Online RL results using verifiable rewards on NaturalReasoning. We apply GRPO to Llama3.1-8B-Instruct using the General Verifier as the reward model.
   Model    MATH    MMLU-Pro    GPQA    Average
   Llama3.1-8B-Instuct    48.74    49.79    29.24    42.59
   GRPO (Step 50)    52.20    49.48    31.47    44.39
These early results suggest that reference answers in NaturalReasoning can be used in RLVR to further enhance reasoning performance, even with limited training. Due to current constraints in computational resources, we leave a more comprehensive exploration of RL-based fine-tuning methods for future work.
Appendix G Evaluating NaturalReasoning Against OpenThoughts
To ensure a fair comparison, we adopt a similar knowledge distillation setup, using a stronger reasoning model (i.e., DeepSeek-R1) as the teacher and Llama-3.1-8B-Instruct as the student. Following the procedure in the OpenThoughts paper [guha2025openthoughtsdatarecipesreasoning], we apply length filtering and then sample 100K questions from each dataset. As shown in Table 8, the model trained on NaturalReasoning outperforms the one trained on OpenThoughts-114k on three out of four benchmarks.
Table 8: Knowledge distillation results comparing training on NaturalReasoning and OpenThoughts. We distill from DeepSeek-R1 into Llama3.1-8B-Instruct using 100K samples from each dataset after applying length filtering following guha2025openthoughtsdatarecipesreasoning.
Train Set GPQA-Diamond MATH-500 MMLU-Pro SuperGPQA
OpenThoughts-114k (100K) 37.8 82.2 59.0 30.0
NaturalReasoning (100K) 43.1 69.0 61.2 32.2
These results show that NaturalReasoning achieves stronger performance on general reasoning and science benchmarks such as GPQA-Diamond, MMLU-Pro, and SuperGPQA [pteam2025supergpqascalingllmevaluation], indicating that it complements existing reasoning datasets like OpenThoughts by providing broader coverage and better generalization.
Appendix H Cross-Domain Generalization of NaturalReasoning
While our main experiments focus on reasoning benchmarks, we also conduct preliminary studies to examine NaturalReasoning ’s applicability to broader knowledge domains. Specifically, we train Llama-3.1-8B-Instruct on 100K randomly sampled NaturalReasoning examples, using DeepSeek-R1 as the teacher model for knowledge distillation. The resulting model is then evaluated on SuperGPQA [pteam2025supergpqascalingllmevaluation], which includes diverse subjects beyond mathematics and science—such as philosophy, law, history, economics, literature, and sociology. The results on non-reasoning categories are shown in Table 9.
Table 9: Cross-domain evaluation results after training on NaturalReasoning. We fine-tune Llama3.1-8B-Instruct on 100K NaturalReasoning examples with DeepSeek-R1 as the teacher model and evaluate on SuperGPQA’s non-reasoning subjects.
Model Philosophy Economics History Education Law Management Literature Sociology
Baseline 29.36 23.17 17.24 26.03 29.63 29.54 23.37 23.78
Train on NR 32.56 34.86 27.49 30.17 29.63 31.14 27.08 33.57
Despite its focus on reasoning-centric tasks, NaturalReasoning improves performance across a wide range of non-reasoning domains. These results highlight its potential as a versatile foundation for training models with broad domain generalization.
Appendix I Prompts